OpenStudy (anonymous):

5(cos^2)60 + 4(sec^2)30 - (tan^2)45 = x Find x

OpenStudy (freckles):

well cos(60) and sec(30) and tan(45) can all be evaluated using the unit circle

OpenStudy (anonymous):

Use the trig table........

OpenStudy (freckles):

do you need help using that chart above ?

OpenStudy (anonymous):

.... I need to the values like cos = 1/2 and all.......

OpenStudy (freckles):

cos(60)=1/2 that is correct so squaring both sides gives cos^2(60)=1/4

OpenStudy (freckles):

sec(30) means you need to find cos(30) and flip whatever number that is

OpenStudy (freckles):

tan(45) means you need to find sin(45) and cos(45) but both sin(45) and cos(45) are the same so tan(45)=1

OpenStudy (freckles):

\[5 \cos^2(60)+4\sec^2(30)-\tan^2(45) \\ 5(\frac{1}{4})+4(\frac{1}{\cos(30)})^2-1\]

OpenStudy (freckles):

so you still need to find cos(30)

OpenStudy (anonymous):

4/3 = sec^2(30)

OpenStudy (freckles):

\[\cos(30)=\frac{\sqrt{3}}{2} \\ \cos^2(30)=\frac{3}{4} \\ \frac{1}{\cos^2(30)}=\frac{4}{3} \\ \sec^2(30)=\frac{4}{3}\] that is right so far you have this now: \[5(\frac{1}{4})+4(\frac{4}{3})-1\]

OpenStudy (freckles):

that can be simplify just be doing a few multiplications and additions(subtractions)

OpenStudy (anonymous):

\[\frac{ 5 }{ 4 } +\frac{ 16 }{ 3 } - 1\]

OpenStudy (freckles):

right so far

OpenStudy (anonymous):

shall I take LCM?

OpenStudy (freckles):

yes which the lcm(4,3,1)=12

OpenStudy (freckles):

\[\frac{5}{4}+\frac{16}{3}-\frac{1}{1} \\ \frac{5(3)}{4(3)}+\frac{16(4)}{3(4)}-\frac{1(12)}{1(12)} \\ \frac{5(3)}{12}+\frac{16(4)}{12}-\frac{12}{12} \\ \frac{5(3)+16(4)-12}{12}\]

OpenStudy (anonymous):

67/12

OpenStudy (freckles):

looks awesome!

OpenStudy (freckles):

and totally correct :)

OpenStudy (anonymous):

Thanks! Please help me a bit more?

OpenStudy (freckles):

Sure. I can take a stab at it.

OpenStudy (triciaal):

@freckles would using identities just take longer?

OpenStudy (freckles):

@triciaal if you want to try something with identities you can

OpenStudy (anonymous):

\[\frac{ cotA-1 }{ 2 -\sec^2(A) } = \frac{ \cot A }{ 1 + tanA }\]

OpenStudy (freckles):

some people don't need to put in terms of sin and cos but I always do because I remember more identities with sin and cos

OpenStudy (anonymous):

Help me?

OpenStudy (freckles):

\[\frac{\frac{\cos(a)}{\sin(a)}-1}{2-\frac{1}{\cos^2(a)}} =\frac{\frac{\cos(a)}{\sin(a)}}{1+\frac{\sin(a)}{\cos(a)}}\] so this is what I would do as a first step second step would be to remove the compound fraction action going on

OpenStudy (anonymous):

I'm suppose to use LHS and get RHS

OpenStudy (freckles):

ok then get rid of the compound fraction action on the the left hand side

OpenStudy (anonymous):

\[\frac{ \cos(a) - \sin(a) }{ \sin(a) }\]

OpenStudy (anonymous):

and for the denominator 2sin^(a)/cos^2(a)

OpenStudy (freckles):

\[\frac{\frac{\cos(a)}{\sin(a)}-1}{2-\frac{1}{\cos^2(a)}} \cdot \frac{\sin(a) \cos^2(a)}{\sin(a)\cos^2(a)} \\ \frac{\cos^3(a)-\cos^2(a)\sin(a)}{2 \cos^2(a)\sin(a)-\sin(a)} \\ \text{ so doing a bit of factoring } \\ \frac{\cos^2(a)(\cos(a)-\sin(a))}{\sin(a)(2 \cos^2(a)-1)} \\ \\ \text{ now recall on the other side we wanted } \cot(a) \\ \text{ on \top } \\ \text{ well we notice } \frac{\cos(a)}{\sin(a)}=\cot(a) \\ \cot(a) \frac{\cos(a)(\cos(a)-\sin(a))}{2 \cos^2(a)-1} \\ \] now somehow we have to show that \[\frac{\cos(a)(\cos(a)-\sin(a))}{2 \cos^2(a)-1}=\frac{1}{1+\tan(a)}\] any ideas @BlackDranzer

OpenStudy (freckles):

if not read this if you want spoilers: \[\frac{\cos(a)(\cos(a)-\sin(a))}{2 \cos^2(a)-1} \cdot \frac{\cos(a)+\sin(a)}{\cos(a)+\sin(a)} \\ \frac{\cos(a)(\cos^2(a)-\sin^2(a))}{(2 \cos^2(a)-1)(\cos(a)+\sin(a))} \\ \text{ well recall double angle identity for } cosine \\ \frac{ \cos(a)}{\cos(a)+\sin(a)}\] now divide both top and bottom by cos(a)

OpenStudy (freckles):

that is I used \[\cos(2a)=\cos^2(a)-\sin^2(a) \\ \text{ and also } \cos(2a)=2\cos^2(a)-1\]

OpenStudy (freckles):

yep and we can bring down that cot(a) factor we had from earlier (though I normally write my fractions like this 1/(1+tan(a)) )

OpenStudy (freckles):

so the key in our proof was the use of double angle identity for cosine

OpenStudy (anonymous):

Nice.. Thanks... I need help in quadratic equations too...

OpenStudy (freckles):

ok let's see it

OpenStudy (freckles):

hey @triciaal did you come up with a way using identities for the first one ?

OpenStudy (freckles):

I honestly haven't given it a try

OpenStudy (anonymous):

An aero plane flying at 1 km horizontally above the ground make an angle of elevation 60 after 10 seconds its 30 find the speed of the aerplane in km/hr

OpenStudy (freckles):

wait what does that mean what is the angle of elevation 60 or 30?

OpenStudy (freckles):

and what is 30?

OpenStudy (anonymous):

from 60 to 30 in secs from the smae point

OpenStudy (freckles):

oh the angle of elevation changed from 60 to 30 in 10 sec gotcha

OpenStudy (freckles):